2.8. Let a, b be elements of a group G. Assume that a has order 5 and a3b = ba3. Prove that ab = ba.


1
Expert's answer
2022-06-22T09:32:41-0400

We have "a^5=e" and "a^3b=ba^3". Applying "a^3" to the second equation we obtain "a^3(a^3b)=a^3(ba^3)=(a^3b)a^3=(ba^3)a^3" and hence "a^6b=ba^6".

As "a^6=a^5\cdot a=ea=a" we obtain "ab=ba".